OpenStudy (wolfshadow009):

MEDAL! I need help solving this. y=x^2 +2x y=3x+20

OpenStudy (mathmusician):

Set x or y equal to 0

OpenStudy (mathmusician):

Doesn't matter which

OpenStudy (lostinmath101):

Combine them or set one variable to 0

OpenStudy (lostinmath101):

Preferably y

OpenStudy (wolfshadow009):

ok

OpenStudy (wolfshadow009):

?

OpenStudy (radar):

y = y x^2 + 2x = 3x + 20 as each is equal to y. Now use algebra. x^2 -x -20 = 0 Factor and solve.

OpenStudy (wolfshadow009):

How do I do that? Like this? 0=3x+20?

OpenStudy (radar):

(x - 5) (x + 4) = 0

OpenStudy (radar):

Solve for x, then get y.

OpenStudy (wolfshadow009):

Ok...I'm still confused though. Sorry

OpenStudy (radar):

Re read my post, it is lined out in a step by step manner.

OpenStudy (radar):

First does, the y = y seem reasonable?

OpenStudy (wolfshadow009):

Yes

OpenStudy (radar):

Given y= x^2 + 2x and y = 3x + 20 if y=y then we must also have: x^2 + 2x = 3x + 20 that has to be true as y=y ..........agree?

OpenStudy (wolfshadow009):

Ohhh yes I see now.

OpenStudy (radar):

Now we isolate by subtracting 3x from both sides giving us: x^2 - x = 20 Now to form a quadratic equation in standard form Subtract 20 from both sides giving us: x^2 - x -20 =0 Have you solved quadratics before?

OpenStudy (wolfshadow009):

Oh ok. And Yes but I never realy got the concept.

OpenStudy (wolfshadow009):

? You just lost me again sorry...

OpenStudy (radar):

y = x^2 + 2x y= 5^2 + 2(5) = 25 + 10 = 35 (5,35) y = -4^2 + 2(-4) = 16 - 8 = 8 (-4, 8) y = 3x + 20 y = 3(5) + 20 = 35 (5,35) y= 3(-4) + 20 = -12 + 20 = 8 (-4,8) the same for both equations two sets of answers

OpenStudy (radar):

I did have an error: Here is the way it should read: The three most common methods are, quadratic formula, completing the square and factoring. Factoring is the way to go on this one, as it easily factors into: (x-5)(x+r) = 0 x - 5 = 0 so that x=5 x + 4 = 0 so that x= -4 Now substitute those values for x and see what y can be. What part of that has caused confusion on your part?

OpenStudy (wolfshadow009):

Where did the 4 and come frome as well as the r?

OpenStudy (wolfshadow009):

and 5*

OpenStudy (radar):

Darn the error is still there: The three most common methods are, quadratic formula, completing the square and factoring. Factoring is the way to go on this one, as it easily factors into: (x-5)(x+4) = 0 x - 5 = 0 so that x=5 x + 4 = 0 so that x= -4 Now substitute those values for x and see what y can be. That fixed, the "r" was supposed to be a 4 lol.

OpenStudy (radar):

The r should of been a 4 so that X^2 - 1x - 20 = 0 is when factored is (x - 5)(x + 4) = 0

OpenStudy (wolfshadow009):

Ohh I get it now.

OpenStudy (radar):